### v2

# Period collapse of Markov triangles

Abstract
Cristofaro-Gardiner and Kleinman [5] showed the complete period collapse of the Ehrhart quasipolynomial of Fibonacci triangles and their irrational limits, by studying the Fourier–Dedekind sums involved in the Ehrhart function of right-angled rational triangles. We generalize this result using integral affine geometrical methods to all Markov triangles, as defined by Vianna [13]. In particular, we show new occurrences of strong period collapse, namely by constructing for each Markov number $p$ a two-sided sequence of rational triangles and two irrational limits with quasipolynomial Ehrhart function of period $p$.
Keywords.
Ehrhart theory, Markov triangles, period collapse, affine integral geometry, symplectic geometry.
Mathematics Subject Classification.
52C05 (53D35)
Acknowledgements.
The problem of studying period collapse phenomena in the Ehrhart theory of Markov triangles was generously proposed by Dan Cristofaro-Gardiner at the Villa Boninchi conference on singular algebraic curves and quantitative symplectic embeddings in September 2025. I thank my doctoral advisor Felix Schlenk for supervising this project as part of my PhD program at the University of Neuchâtel. I also thank Joel Schmitz for his explanations of integral affine geometry, and Nadine Fares and Lama Yaacoub for their helpful comments.
1 Introduction and main results
Outline of the paper.
We begin by recalling some elements of integral affine geometry in §1.1 and Ehrhart theory in §1.2, and then define Markov triangles in §1.3. The main results are Theorems 1.25 and 1.26, stated in §1.4. In §1.5 we discuss connections of the Ehrhart theory of our triangles to symplectic embedding problems. All proofs are in §2.
1.1 Integral affine geometry
Basic concepts.
Here we define the notions of planar integral affine geometry used in this paper. We refer the interested reader to [10] for an extensive account of the theory. By a polygon $P$, we mean a two-dimensional convex polytope, i.e. the convex hull of finitely many points in $\mathbb{R}^{2}$, called the vertices of $P$, while the boundary line segments connecting vertices are called the edges of $P$. A polygon is said to be rational resp. integral if its vertices have rational resp. integral coordinates.
For rational polygons we can define some additional notions: consider three distinct points $p_{1},p_{2},p_{3}\in\mathbb{R}^{2}$ with rational coordinates.
  • The primitive integral vector $v_{1,2}$ along the line segment $p_{1},p_{2}]$ is defined to be the unique vector with coprime integer coordinates obtained by multiplying $p_{2}-p_{1}$ by $k\in\mathbb{R}_{>0}$. The affine length of $p_{1},p_{2}]$ is then $r\vcentcolon=\frac{1}{k}$, namely the unique $r\in\mathbb{R}_{>0}$ such that $p_{2}-p_{1}=rv_{1,2}$.
  • The determinant of the angle at $p_{1}$ is defined as $|\det(v_{1,2},v_{1,3})|$, namely the absolute value of the determinant of the matrix formed by the primitive integral vectors $v_{1,2}$ and $v_{1,3}$ along the segments $p_{1},p_{2}]$ and $p_{1},p_{3}]$, respectively. The integral bisector of the angle at $p_{1}$ is the affine line through $p_{1}$ spanned by $v_{1,2}+v_{1,3}$.
An affine transformation $\varphi\in\mathbb{R}^{2}\rtimes\operatorname{GL}(2,\mathbb{R})$ is said to be integral if it preserves the lattice $\mathbb{Z}^{2}$, i.e. $\varphi(\mathbb{Z}^{2})=\mathbb{Z}^{2}$. The group of integral affine transformations is the semidirect product $\mathbb{Z}^{2}\rtimes\operatorname{GL}(2,\mathbb{Z})$ of the group of translations along integral vectors with the group of integral linear transformations. Two polygons $P,P^{\prime}$ are said to be integrally congruent if there exists an element $\varphi\in\mathbb{Z}^{2}\rtimes\operatorname{GL}(2,\mathbb{Z})$ taking $P$ to $P^{\prime}$.
Figure 1: Examples of integrally congruent triangles
Half-shears.
Given a vector $v\in\mathbb{R}^{2}$, we denote by $H^{+}_{v}$ the half-plane lying to the right of the line spanned by $v$, and by $H^{-}_{v}$ the one lying to its left; more technically, $H_{v}^{+}$ resp. $H_{v}^{-}$ is the set of vectors $u\in\mathbb{R}^{2}$ for which $\det(u,v)\geq 0$ resp. $\det(u,v)\leq 0$.
Definition 1.1.
Given an integral vector $v=(v_{1},v_{2})\in\mathbb{Z}^{2}$, we define the shear with respect to $v$ as integral linear map $\varphi_{v}\in\operatorname{GL}(2,\mathbb{Z})$ given by $u\mapsto u+\det(v,u)v$, or in matrix form
$\varphi_{v}=\begin{pmatrix}1-v_{1}v_{2}&v_{1}^{2}\\ -v_{2}^{2}&1+v_{1}v_{2}\end{pmatrix}$.
We also define the half-shear with respect to $v$ as the map $\varphi^{1/2}_{v}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ with
$\varphi^{1/2}_{v}(u)\vcentcolon=\begin{cases}\varphi_{v}(u)&\text{for }u\in H_{v}^{+}\\ u&\text{otherwise}\end{cases}.$
Note that $\varphi^{1/2}_{v}$ is well-defined, since $\varphi_{v}$ fixes the line spanned by $v$.
In §1.3 we will need the following conjugated half-shears.
Definition 1.2.
Let $P\subset\mathbb{R}^{2}$ be a rational polygon, $p\in P$ a vertex of $P$ and $v_{p}$ the primitive integral vector along the integral bisector of the angle at $p$. The geometric mutation $M_{p}\vcentcolon=M_{p,P}$ of $P$ at $p$ is defined as the map obtained by conjugating the half-shear with respect to $v_{p}$ by translation along $p$, i.e.
$M_{p}\vcentcolon=T_{p}\circ\varphi^{1/2}_{v_{p}}\circ T_{-p}.$
Remark 1.3.
The geometric mutation of a rational polygon at an integral vertex preserves the integer lattice. In particular, half-shears preserve the integer lattice.
1.2 Ehrhart theory
Basic notions.
We refer to [3, Chapter 3] for an excellent exposition of the topic. Let $P\subset\mathbb{R}^{2}$ be a polygon.
Definition 1.4.
The Ehrhart function of $P$ is the function in $t\in\mathbb{N}\cup\{0\}$ defined as the lattice-point enumerator of $t$-dilates of $P$; it is denoted by $L_{P}$, where
$L_{P}(t)\vcentcolon=\#(tP)\cap\mathbb{Z}^{2}.$
Two polygons are said to be Ehrhart equivalent if they have identical Ehrhart functions. We observe that half-shears preserve Ehrhart functions:
Proposition 1.5.
Let $v$ be an integral vector, and consider the half-shear $\varphi^{1/2}_{v}$ along $v$. Then $\varphi^{1/2}_{v}(P)$ and $P$ are Ehrhart equivalent.
The first fundamental result of Ehrhart theory is due to Ehrhart himself [7].
Theorem 1.6 (Ehrhart).
If $P$ is integral, then $L_{P}$ is a polynomial in $t$.
If $P$ is rational, we define its denominator $\operatorname{den}(P)$ to be the least common multiple of the denominators of the coordinates of all vertices of $P$ (see Figure 2). The $\operatorname{den}(P)$-dilate of $P$ is integral, and hence its Ehrhart function is a polynomial.
[…]
Figure 2: Example of a triangle with denominator $24$
Definition 1.7.
A function $f:\mathbb{N}\cup\{0\}\rightarrow\mathbb{R}$ is said to be quasipolynomial if it is periodically polynomial, i.e. there exists $T\in\mathbb{Z}_{>0}$ and a collection of $T$ polynomials $\{f_{k}\}_{k=0}^{T-1}$ such that for all $t\in\mathbb{N}\cup\{0\}$, $f(t)=f_{k}(t)$ whenever $t\equiv k$ mod $T$. We call $T$ the period of $f$.
Theorem 1.8.
If $P$ is rational, then $L_{P}$ is quasipolynomial with period $T$ dividing $\operatorname{den}(P)$.
Period collapse.
Given a rational polygon $P$, we say that period collapse occurs if the period $T$ of $L_{P}$ is strictly smaller than $\operatorname{den}(P)$. Period collapse is said to be complete if the period equals $1$, i.e. $L_{P}$ is polynomial, in which case we also say that $P$ is pseudo-integral in accordance with the terminology used by Cristofaro-Gardiner et al. [6]. Note that while period collapse is not the behaviour that would be expected a priori, it is not a rare phenomenon: McAllister and Woods [11] showed that for an arbitrary integer $D$ and any divisor $d$ of $D$, there exists a rational triangle with denominator $D$ and Ehrhart quasipolynomial with period $d$. We may also extend the discussion to include irrational polygons; we say that an irrational polygon $P$ is pseudo-rational if $L_{P}$ is a quasipolynomial. Note that this is an extreme form of period collapse if we view $P$ as a rational polygon with infinite denominator.
Example 1.9.
The Fibonacci sequence $(\operatorname{Fib}_{n})_{n\in\mathbb{N}}$ is defined recursively by
$\operatorname{Fib}_{1}=\operatorname{Fib}_{2}=1,\,\operatorname{Fib}_{n+2}=\operatorname{Fib}_{n+1}+\operatorname{Fib}_{n}.$
We construct a family of triangles corresponding to the odd-indexed Fibonacci numbers as follows: given a pair $(r,s)$ of real numbers, the $(r,s)$-triangle $\mathcal{F}_{(r,s)}$ is defined by its vertices $(0,0)$, $(r,0)$, and $(0,s)$. The $n$-th Fibonacci triangle is defined as the $(g_{n}\vcentcolon=\frac{\operatorname{Fib}_{2n-1}}{\operatorname{Fib}_{2n-3}},\frac{\operatorname{Fib}_{2n-3}}{\operatorname{Fib}_{2n-1}})$-triangle, for $n\geq 2$ [^2] We define the first Fibonacci triangle to be the standard $2$-dimensional simplex $\mathcal{F}_{(1,1)}$.. Recall that the sequence $\operatorname{Fib}_{n+1}/\operatorname{Fib}_{n}$ converges to the golden ratio $\tau=\frac{1+\sqrt{5}}{2}$ as $n\rightarrow\infty$, which allows us to define the limiting triangle $\mathcal{F}_{\infty}\vcentcolon=\lim_{n\rightarrow\infty}\mathcal{F}_{(g_{n},1/g_{n})}=\mathcal{F}_{(\tau^{2},1/\tau^{2})}$ (see Figure 3). $\mathcal{F}_{\infty}$ is irrational. Cristofaro-Gardiner and Kleinman [5] showed that complete period collapse occurs for all Fibonacci triangles as well as their limit $\mathcal{F}_{\infty}$.
[…]
Figure 3: The limiting Fibonacci triangle
The goal of this paper is to generalize the period collapse phenomenon for Fibonacci triangles to a larger class of triangles, the so-called Markov triangles.
1.3 Markov triangles
The Markov tree.
The Diophantine equation
$p_{1}^{2}+p_{2}^{2}+p_{3}^{2}=3p_{1}p_{2}p_{3}$ (1)
is called Markov’s equation, and a solution $(p_{1},p_{2},p_{3})\in\mathbb{Z}_{\geq 1}^{3}$ of (1) is called a Markov triple, with elements called Markov numbers. We will need the following fact from [2, Corollary 3.4].
Fact 1.10.
The elements of a Markov triple are pairwise coprime.
[…]
(1,1,1)(1,1,1)$(1,2,1)$ $\underline{\overline{(1,5,2)}}$ $\overline{(1,13,5)}$ $(1,34,13)$ $\overline{(5,194,13)}$ $\underline{(2,29,5)}$ $\underline{(5,433,29)}$ $(2,169,29)$
Figure 4: The first five generations of the Markov tree
Let $p\notin\{1,2\}$ be a Markov number and $\mathfrak{m}$ a Markov triple in which $p$ is the largest number. We denote by $\Lambda(\mathfrak{m})$ the subgraph of the Markov tree consisting of all triples obtained by an iteration of mutations of $\mathfrak{m}$ preserving $p$. $\Lambda(\mathfrak{m})$ is a bivalent subtree rooted in $\mathfrak{m}=\vcentcolon(a,p,b)$ consisting of two branches: the $B_{p}^{a}$-branch obtained by mutating $(a,p,b)$ first at $b$, and the $B_{p}^{b}$-branch obtained by mutating $(a,p,b)$ first at $a$, each of which then extends by mutations that preserve $p$ and the largest element of the triple. As an example, the subtree formed by the vertices which are either underlined or overlined in Figure 4 is $\Lambda(1,5,2)$. The first two vertices of the branch $B_{5}^{1}$ are the overlined triples and those of the branch $B_{5}^{2}$ are the underlined triples.
In the case $p=1$, since we omitted two of the three identical subtrees adjacent to $(1,1,1)$, the subgraph rooted in $(1,1,1)$ obtained via mutations preserving $1$ is a branch, with a particularity: given a Markov triple $(1,p_{2},p_{3})$, we have $\widehat{p_{2}}=3p_{3}-p_{2}$, which is the recursive equation defining odd-indexed Fibonacci numbers,i.e. $\operatorname{Fib}_{2n+3}=3\operatorname{Fib}_{2n+1}-\operatorname{Fib}_{2n-1}$. Since $\operatorname{Fib}_{1}=1$ and $\operatorname{Fib}_{3}=2$, it follows by induction that the triples of this branch are given by the odd-indexed Fibonacci numbers, i.e the $n$-th triple (for $n\geq 2$) is
$(1,\operatorname{Fib}_{2n-1},\operatorname{Fib}_{2n-3}).$
Accordingly, we call this branch the Fibonacci branch of the Markov tree. A similar situation is observed in the case $p=2$, where the subgraph obtained by mutations preserving $2$ and rooted in $(1,2,1)$ is a branch with $n$-th triple (for $n\geq 2$)
$(1,\operatorname{Pell}_{2n-1},\operatorname{Pell}_{2n-3})$
where $(\operatorname{Pell}_{n})_{n\in\mathbb{N}}$ is the Pell sequence, defined recursively by
$\operatorname{Pell}_{0}=0,\,\operatorname{Pell}_{1}=1,\,\operatorname{Pell}_{n+2}=2\operatorname{Pell}_{n+1}+\operatorname{Pell}_{n}.$
Accordingly, we call this branch the Pell branch of the Markov tree. Note that the Fibonacci and Pell branches are the peripheral branches of the tree. Many more interesting results and open problems on Markov numbers are surveyed in [2].
Markov triangles.
Motivated by the complete period collapse of Fibonacci triangles, we study a larger class of rational triangles representing all Markov triples defined by Vianna [13] and Evans [8, Appendix I].
Definition 1.11.
Let $(p_{1},p_{2},p_{3})$ be a Markov triple. A Markov triangle $\Delta_{(p_{1},p_{2},p_{3})}$ representing $(p_{1},p_{2},p_{3})$ is defined recursively as follows:
  • $\Delta_{(1,1,1)}$ is any triangle with edges having affine length $1$ and vertices having angles of determinant $1$.
  • For the $n+1$-st generation Markov triple $(p_{1},\widehat{p_{2}},p_{3})$ obtained by mutating the $n$-th generation triple $(p_{1},p_{2},p_{3})$ at $p_{2}$, $\Delta_{(p_{1},\widehat{p_{2}},p_{3})}$ is defined as the image of the geometric mutation of $\Delta_{(p_{1},p_{2},p_{3})}$ at the $p_{2}$-vertex, namely
    $\Delta_{(p_{1},\widehat{p_{2}},p_{3})}\vcentcolon=M_{p_{2}}(\Delta_{(p_{1},p_{2},p_{3})})$
    where $M_{p_{2}}\vcentcolon=M_{p_{2},\Delta_{(p_{1},p_{2},p_{3})}}$ is given in Definition 1.2.
$1M_{1}2255$$29$
Figure 5: Geometric mutation of a $(1,5,2)$-triangle at the $1$-vertex
We say that two triangles $\Delta$ and $\Delta^{\prime}$ are Markov equivalent if they represent the same Markov triple.
Remark 1.12.
It is not true in general that the image of a triangle $\Delta$ under geometric mutation $M_{p}$ at a vertex $p$ is again a triangle: it may be a quadrilateral. However in the case of Markov triangles, $M_{p}(\Delta)$ is a triangle, as shown by Evans [8, Theorem 8.21]. We also give a computational proof of this fact in §2.3.
Remark 1.13.
The equivalence classes of Markov triangles are invariantunder the $\mathbb{R}^{2}\rtimes\operatorname{GL}(2,\mathbb{Z})$-action, on the one hand because geometric mutations commute with the $\operatorname{GL}(2,\mathbb{Z})$-action, and on the other hand because the class of Markov triangles representing $(1,1,1)$ is by definition invariant under $\mathbb{R}^{2}$-translations.
It will be useful for our purposes to fix a model triangle in each Markov class.
Definition 1.14.
Let $\mathfrak{m}\vcentcolon=(p_{1},p_{2},p_{3})$ be a Markov triple. The companion numbers $q_{1}^{\pm}\in\mathbb{Z}_{p_{1}}$ of $p_{1}$ with respect to $\mathfrak{m}$ are $q_{1}^{\pm}\vcentcolon\equiv\pm 3p_{3}p_{2}^{-1}$ mod $p_{1}$.
Remark 1.15.
The set $\{q_{1}^{\pm}\}$ does not depend on the order of the triple $\mathfrak{m}$, since the congruence $p_{2}^{2}+p_{3}^{2}\equiv 0$ mod $p_{1}$ implies that $3p_{2}p_{3}^{-1}\equiv-3p_{3}p_{2}^{-1}$ mod $p_{1}$.
We will show in Proposition 2.7 that every Markov triangle $\Delta$ representing a triple $(p_{1},p_{2},p_{3})$ is $\mathbb{R}^{2}\rtimes\operatorname{GL}(2,\mathbb{Z})$-congruent to the triangle $\Delta(q_{1})$ with $p_{1}$-, $p_{3}$-, and $p_{2}$-vertices
$\Bigg\{(0,0),\Big(\frac{(p_{1}q_{1}-1)p_{2}}{p_{1}p_{3}},\frac{p_{1}p_{2}}{p_{3}}\Big),\Big(\frac{p_{3}}{p_{1}p_{2}},0\Big)\Bigg\}$ (2)
for any companion number $q_{1}\vcentcolon\equiv q_{1}^{+}\equiv 3p_{3}p_{2}^{-1}$ mod $p_{1}$ of $p_{1}$.
Definition 1.16.
A Markov triangle representing the triple $(p_{1},p_{2},p_{3})$ with coordinates given by (2) is said to be in a standard $p_{1}$-position.
This definition allows us to view the Fibonacci triangles from Example 1.9 as Markov triangles representing the triples $(1,\operatorname{Fib}_{2n-1},\operatorname{Fib}_{2n+1})$ in standard $1$-position.
Remark 1.17.
The $\operatorname{GL}(2,\mathbb{Z})$-class of $\Delta(q_{1})$ depends only on the congruence class of $q_{1}$ modulo $p_{1}$: for $q_{1}^{\prime}\in\mathbb{Z}$ with $q_{1}-q_{1}^{\prime}=kp_{1}$, $\Delta(q_{1}^{\prime})$ is the image of $\Delta(q_{1})$ under
$\begin{pmatrix}1&-k\\ 0&1\end{pmatrix}$.
Moreover, the $\operatorname{GL}(2,\mathbb{Z})$-class of $\Delta(q_{1})$ does not depend on the order of $(p_{1},p_{2},p_{3})$: the triangle $\Delta^{\prime}$ obtained by swapping $p_{2}$ and $p_{3}$ in the definitions of $\Delta(q_{1})$ and of $q_{1}$ is the image of $\Delta(q_{1})$ under
$\psi_{p_{1}}^{p_{2},p_{3}}\vcentcolon=\begin{pmatrix}-1-q_{1}p_{1}&q_{1}^{2}\\ p_{1}^{2}&1-q_{1}p_{1}\end{pmatrix}$.
1.4 Ehrhart theory of Markov triangles
In the following, we define the integral barycentre of a triangle to be the intersection of its integral bisectors, which is well-defined by Lemma 2.3. Recall that the Ehrhart function of a triangle $\Delta$ does not change under a half-shear of $\Delta$, but might change under an $\mathbb{R}^{2}$-translation. Given a Markov triple $(p_{1},p_{2},p_{3})$, we consider:
  • Markov triangles representing $(p_{1},p_{2},p_{3})$ in a standard $p_{1}$-position,
    xxyy22552929
    Figure 6: A $(2,5,29)$-triangle in standard $2$-position
  • Markov triangles representing $(p_{1},p_{2},p_{3})$ in a barycentric position, i.e. with the integral barycentre at the origin of the $\mathbb{R}^{2}$-plane. In particular, such a Markov triangle $\Delta$ obtained by translating a triangle representing $(p_{1},p_{2},p_{3})$ in standard $p_{1}$-position (so that $\Delta$ has an edge parallel to the $x$-axis) is said to be in a standard $p_{1}$-barycentric position. [^4] The triangles depicted in Figure 7 are only illustrative; typical Markov triangles in barycentric position are very thin and have one vertex extremely close to the origin, and hence their depiction would not help visualization.
    xxyy22552929
    xxyy29295522
    Figure 7: $(2,5,29)$-triangles in a generic barycentric position (left) and in a standard $2$-barycentric position (right)
Remark 1.18.
Since every $(p_{1},p_{2},p_{3})$-triangle in barycentric position is $\operatorname{GL}(2,\mathbb{Z})$-congruent to one in standard $p_{1}$-barycentric position, it will suffice for us to consider the latter position.
Remark 1.19.
More precisely, a triangle $\Delta$ representing the triple $(p_{1},p_{2},p_{3})$ in standard $p_{1}$-barycentric position is obtained from a triangle $\Delta(q_{1})$ in standard $p_{1}$-position by translation along the vector pointing from the integral barycentre $\beta^{(p_{1})}$ of $\Delta(q_{1})$ to the origin. By the computation done in the proof of Proposition 1.21, we have $\beta^{(p_{1})}=(\frac{q_{1}}{3p_{1}},\frac{1}{3})$; hence $\Delta=T_{(\frac{-q_{1}}{3p_{1}},\frac{-1}{3})}(\Delta(q_{1}))$.
The rational case.
The following result shows that the Ehrhart theory of Markov triangles in standard $p_{1}$-barycentric position is particularly simple.
Proposition 1.20.
All Markov triangles in standard $p_{1}$-barycentric position are Ehrhart equivalent, having Ehrhart quasipolynomial with period $3$.
We note that the denominator of a Markov triangle representing $(p_{1},p_{2},p_{3})$ in standard $p_{1}$-position is $p_{1}p_{2}p_{3}$. The following result says that in this case there is strong, though not complete, period collapse.
Proposition 1.21.
The Ehrhart quasipolynomial of a $(p_{1},p_{2},p_{3})$-triangle in standard $p_{1}$-position has period a divisor of $p_{1}$.
The complete period collapse of the Fibonacci triangles proved in [5] is therefore a direct consequence of Proposition 1.21. Another direct implication is to the Markov triangles representing the triples lying on the $2$-branch, which we call the Pell triangles.
Corollary 1.22.
The Ehrhart quasipolynomial of a Pell triangle in standard $2$-position has period $2$.
The limiting case.
Take some Markov number $a$ as the largest element of a Markov triple $(b,a,c)$, and consider the bivalent subtree $\Lambda(b,a,c)$. We study the evolution of the representative Markov triangles while going down the $B_{c}$-branch of $\Lambda(b,a,c)$: fix a companion number $q\equiv 3cb^{-1}$ mod $a$ of $a$ with respect to $(a,b,c)$, and consider the triple $(a,b_{n},c_{n})$ obtained by first mutating $(a,b,c)$ at $b$ and then applying n1n-1 increasing mutations preserving $a$, with $b_{0}\vcentcolon=b$, $c_{0}\vcentcolon=c$, and $b_{n+1}\vcentcolon=c_{n}$, $c_{n+1}\vcentcolon=3ac_{n}-b_{n}$. Let $\Delta^{a}_{n}\vcentcolon=\Delta_{n}^{a}(q,c)$ be the Markov triangle representing $(a,b_{n},c_{n})$ in standard $a$-position with the $b_{n}$-vertex on the $x$-axis, and let $\Delta^{a,\beta}_{n}\vcentcolon=\Delta^{a,\beta}_{n}(q,c)$ be the Markov triangle representing $(a,b_{n},c_{n})$ in standard $a$-barycentric position obtained by a translation of $\Delta_{n}^{a}$.
Remark 1.23.
Going down the $B_{b}$-branch of the bivalent tree instead of the $B_{c}$- branch yields a second pair of sequences of representative Markov triangles $\Delta_{n}^{a}(-q,b)$ and $\Delta_{n}^{a,\beta}(-q,b)$, both of which are constructed simply by swapping the letters $b$ and $c$ in the definitions of their $B_{c}$-branch homologues. For the sake of clarity we will restrict our attention in the rest of the paper to the sequences $\Delta_{n}^{a}(q,c)$ and $\Delta_{n}^{a,\beta}(q,c)$, but note that the same results hold for the sequences $\Delta_{n}^{a,\beta}(-q,b)$ and $\Delta_{n}^{a,\beta}(-q,b)$.
Let us first investigate the sequence $(\Delta_{n}^{a})$ of triangles in standard $a$-position; we note that $\Delta_{n+1}^{a}$ is obtained by geometrically mutating $\Delta_{n}^{a}$ at the $b_{n}$-vertex for all $n\geq 2$. The primitive vectors $v_{a}^{(1)}\vcentcolon=v_{a,b_{n}}=(1,0)$ and $v_{a}^{(2)}\vcentcolon=v_{a,c_{n}}=(aq-1,a^{2})$ are preserved by the geometric mutations involved in the definition of $\Delta_{n}^{a}$, and hence $\Delta_{n}^{a}$ is entirely defined by the affine lengths of the $a,b_{n}]-$ and $a,c_{n}]-$edges, namely $\frac{c_{n}}{ab_{n}}$ and $\frac{b_{n}}{ac_{n}}$ respectively, as illustrated in Figure 8. We note that the ratio $\frac{ab_{n}}{c_{n}}$ has a nice limiting behaviour,
$\lim_{n\rightarrow\infty}\frac{ab_{n}}{c_{n}}=\frac{1}{\lambda(a)}$
where $\lambda(a)\vcentcolon=\frac{3+\sqrt{9-4/a^{2}}}{2}$ is the Lagrange number corresponding to $a$, see [4, Section 3] [^5] Brendel and Schlenk define $\lambda(a)$ as $\sqrt{9-4/a^{2}}$. We use an alternative definition in the present paper for the sake of clarity.. We may then compute the limits of the affine lengths of the $a,b_{n}]-$ and $a,c_{n}]-$edges
$\lim_{n\rightarrow\infty}\frac{c_{n}}{ab_{n}}=\lambda(a)\qquad\text{and}\qquad\lim_{n\rightarrow\infty}\frac{b_{n}}{ac_{n}}=\frac{1}{a^{2}\lambda(a)}$ (3)
which allows us to define the triangle $\Delta_{\infty}^{a}$ with $a$-vertex at the origin, $b_{\infty}$-vertex at an affine distance $\lambda(a)$ from $a$ along $v_{a}^{(1)}$, and $c_{\infty}$-vertex at an affine distance $\frac{1}{a^{2}\lambda(a)}$ from $a$ along $v_{a}^{(2)}$. $\Delta_{\infty}^{a}$ is in fact the limit of the Markov triangle sequence in question.
[…]
Figure 8: Exaggerated representation of a sequence of Markov triangles in standard $a$-position
Proposition 1.24.
$(\Delta_{n}^{a})$ Hausdorff converges to $\Delta_{\infty}^{a}$.
By Proposition 1.21, the Ehrhart function of $\Delta_{n}^{a}$ is a quasipolynomial with period a divisor of $a$, for all $n\in\mathbb{N}$. It turns out that this period collapse also occurs in the limiting case:
Theorem 1.25.
The $a$-dilation of the limiting irrational triangle $\Delta_{\infty}^{a}$ is Ehrhart equivalent to the $a$-dilation of $\Delta_{n}^{a}$ for any $n\in\mathbb{N}$. In particular, $\Delta_{\infty}^{a}$ is pseudo-rational with period $a$.
This period collapse is not a priori a result of Hausdorff convergence, as pointed out in Remark 2.12. Let us now consider the sequence $(\Delta_{n}^{a,\beta})$ of Markov triangles in standard $a$-barycentric position. It follows clearly from Proposition 1.24 and the fact that $\Delta_{n}^{a,\beta}$ is the $(\frac{-a}{3q},\frac{-1}{3})$-translation of $\Delta_{n}^{a}$ by Remark 1.19, that $(\Delta_{n}^{a,\beta})$ Hausdorff converges to the triangle $\Delta_{\infty}^{a,\beta}$ defined as the $(\frac{-a}{3q},\frac{-1}{3})$-translation of $\Delta_{\infty}^{a}$.
Theorem 1.26.
The limiting irrational triangle $\Delta_{\infty}^{a,\beta}$ in standard $a$-barycentric position is Ehrhart equivalent to $\Delta_{n}^{a,\beta}$ for any $n\in\mathbb{N}$. In particular, $3\Delta_{\infty}^{a,\beta}$ is pseudo-integral.
Open problem.
Recall from Example 1.9 the triangle $\mathcal{F}_{(b/c,c/b)}$ for integers $b,c\in\mathbb{N}$ with one vertex at the origin and the two other vertices at $\frac{b}{c}(0,1)$ resp. $\frac{c}{b}(1,0)$. Cristofaro-Gardiner and Kleinman show in [5, Theorem 1.6] that $\mathcal{F}_{(b/c,c/b)}$ is pseudo-integral if and only if $b$ and $c$ are consecutive odd-indexed Fibonacci numbers. This motivates the following problem: take coprime integers $a,q\in\mathbb{N}$. For integers $b,c\in\mathbb{N}$, define the triangle $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ with one vertex at the origin and the two other vertices at $\frac{b}{c}(aq-1,a^{2})$ resp. $\frac{c}{b}(1,0)$. Proposition 1.21 shows that $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ is pseudo-integral if $(a,b,c)$ is a Markov triple and $q\equiv 3cb^{-1}$ mod $a$, since in this case $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ is the $a$-dilate of a Markov triangle representing $(a,b,c)$ in standard $a$-position. This generalizes one direction of [5, Theorem 1.6]. Is the converse true, i.e. do there exist pseudo-integral triangles $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ which are not Markov? A similar question may be asked in the barycentric setting: Proposition 1.20 says that the $\frac{3}{a}$-dilate of the $(\frac{-q}{3a},\frac{-1}{3})$-translate of the triangle $\mathcal{F}^{(p,q)}_{(b/c,c/b)}$ is pseudo-integral if $(a,b,c)$ is a Markov triple and $q\equiv 3cb^{-1}$ mod $a$. Is the converse true?
1.5 Connections to symplectic geometry
The rational triangles whose Ehrhart theory we study in this paper are special in two ways: they admit complete or at least strong period collapse, and their irrational limit triangles also have this property. This Ehrhart-peculiarity may be a shadow of their relevance in symplectic geometry.
Indeed, our rational triangles arise as bases of almost toric fibrations of the complex projective plane $\mathbb{CP}^{2}$, see [13], [8]. Furthermore, these triangles are the moment map images of exactly those weighted projective planes that admit smoothings to $\mathbb{CP}^{2}$, see [9]. On the other hand, these triangles also describe certain rational homology ellipsoids, that in the special case of Fibonacci triangles are usual ellipsoids. As a consequence, for every Markov number, the problem of symplectically embedding rational homology ellipsoids into a minimal $\mathbb{CP}^{2}$ leads to an infinite staircase, see [1].
For Fibonacci triangles, this relation between Ehrhart theory and symplectic embedding problems was noticed and studied earlier in [5] and [6], and the corresponding problem of symplectically embedding usual ellipsoids into $\mathbb{CP}^{2}$, but also into the 4-ball, led to the Fibonacci staircase in [12].
2 Proofs
2.1 Integral barycentre of a rational triangle
Let $p,p_{0}\in\mathbb{R}^{2}$ be rational vectors, $v\in\mathbb{Z}^{2}$ a primitive integral vector. We define the affine distance from $p$ to the affine line $\ell\vcentcolon=\ell(p_{0},v)$ through $p_{0}$ spanned by $v$ to be the absolute value of the determinant of $p-p_{0}$ with $v$, i.e. $|\det(p-p_{0},v)|$.
Remark 2.1.
The affine distance from $p$ to $\ell$ does not depend on the choice of $p_{0}$: given another point $p_{0}^{\prime}\in\ell$, we have
$\det(p-p_{0}^{\prime},v)=\det(p-p_{0}+p_{0}-p_{0}^{\prime},v)=\det(p-p_{0},v)$
since $p_{0}-p_{0}^{\prime}$ is a scalar multiple of $v$.
In classical geometry, the bisector of an angle $\alpha$ in the plane is defined as the set of all points lying at an equal distance from both lines spanning $\alpha$. The following result asserts that a similar definition holds true for integral bisectors.
Lemma 2.2.
Let $u,v\in\mathbb{Z}^{2}$ be linearly independent primitive integral vectors. A vector $w\in\mathbb{R}^{2}$ is a scalar multiple of $u+v$ if and only if $\det(w,u)=\det(v,w)$.
Proof.
Assume first that $w=\lambda(u+v)$ for some $\lambda\in\mathbb{R}$. Then
$\det(w,u)=\det(\lambda(u+v),u)=\lambda\det(v,u)=\det(v,\lambda(u+v))=\det(v,w).$
Conversely assume that $w$ is not a multiple of $u+v$. Since $u$ and $v$ span $\mathbb{R}^{2}$, then we may write $w=\lambda_{1}u+\lambda_{2}v$ for some scalars $\lambda_{1}\neq\lambda_{2}$. Then
$\det(w,u)=\lambda_{2}\det(v,u)\neq\lambda_{1}\det(v,u)=\det(v,w).\qed$
Let $\Delta\subset\mathbb{R}^{2}$ be a rational triangle with vertices $p_{1},p_{2},p_{3}$. Unwinding the definitions, Lemma 2.2 says that the integral bisector of the angle at $p_{1}$ is the set of points at equal affine distance from the $p_{1},p_{2}]-$ and the $p_{1},p_{3}]-$edges.
Lemma 2.3.
The three integral bisectors of $\Delta$ intersect at a single point.
Proof.
Let $\ell_{i}$ be the integral bisector of the angle at $p_{i}$ (for $i\in\mathbb{Z}_{3}$), and let $\beta$ be the intersection of $\ell_{1}$ and $\ell_{2}$. Then we have
$|\det(\beta-p_{3},v_{1,3})|$ $=|\det(\beta-p_{1},v_{1,3})|=|\det(v_{1,2},\beta-p_{1})|$
$=|\det(v_{1,2},\beta-p_{2})|=|\det(\beta-p_{2},v_{2,3})|=|\det(\beta-p_{3},v_{2,3})|$
where the first, third and fifth equalities follow from Remark 2.1, while the second and fourth equalities follow from the definition of $\beta$ and Lemma 2.2. We deduce, also by Lemma 2.2, that $\beta-p_{3}$ is a scalar multiple of $v_{3,1}+v_{3,2}$, i.e. it lies on $\ell_{3}$.
$p_{1}$$p_{2}$$p_{3}$$v_{1,3}$$v_{3,1}$$v_{1,2}$$v_{2,1}$$v_{3,2}$$v_{2,3}$$\ell_{1}$$\ell_{3}$$\ell_{2}$$\beta$
Figure 9: Proof of Lemma 2.3
As mentioned in §1.4, we call the intersection of the integral bisectors of $\Delta$, the integral barycentre of $\Delta$.
Remark 2.4.
The integral barycentre of $\Delta$ is equivalently defined as the point lying at an equal affine distance from the three edges of $\Delta$.
2.2 On model Markov triangles
We introduce an alternative version of Markov triangles, which turns out to be equivalent to Definition 1.11.
Definition 2.5.
Let $(p_{1},p_{2},p_{3})$ be a Markov triple. We define a pseudo-Markov triangle $\Delta_{(p_{1},p_{2},p_{3})}$ representing $(p_{1},p_{2},p_{3})$ to be any triangle (with vertices denoted by $p_{1},p_{2},p_{3}$) satisfying the following conditions
  • the edge opposite to $p_{i}$ has affine length $\frac{p_{i}}{p_{i+1}p_{i+2}}$ (where $i\in\mathbb{Z}_{3}$),
  • the determinant of the angle at the $p_{i}$-vertex is $p_{i}^{2}$.
To prove the equivalence of Definitions 1.11 and 2.5, we need the following two facts about companion numbers of Markov numbers.
Lemma 2.6.
Consider a Markov triple $(p_{1},p_{2},p_{3})$ and take a companion number $q_{1}$ of $p_{1}$ with respect to $(p_{1},p_{2},p_{3})$. Then $q_{1}$ and $p_{1}$ are coprime. Moreover, $q_{1}$ is also a companion of $p_{1}$ with respect to $(p_{1},\widehat{p_{2}},p_{3})$.
Proof.
First we show that $p_{1}$ is relatively prime to $3$: assume by way of contradiction that $3\mid p_{1}$, so that the Markov equation implies $p_{2}^{2}+p_{3}^{2}\equiv 0$ mod $3$. Since $x^{2}\equiv 0$ or $x^{2}\equiv 1$ mod $3$ for all $x$, it follows that $p_{2}^{2}\equiv p_{3}^{2}\equiv 0$ mod $3$ and hence $3$ divides all three elements of the triple $(p_{1},p_{2},p_{3})$, contradicting their pairwise coprimality.Thus $p_{1}$ is coprime to any number congruent to $\pm 3p_{3}p_{2}^{-1}$ mod $p_{1}$, in particular $q_{1}$. In addition, we have
$q_{1}\equiv\pm 3p_{2}p_{3}^{-1}$ $\equiv\mp 3(-p_{2})p_{3}^{-1}$ $\text{mod }p_{1}$
$\equiv\mp 3(3p_{1}p_{3}-p_{2})p_{3}^{-1}$ $\text{mod }p_{1}$
$\equiv\mp 3\widehat{p_{2}}p_{3}^{-1}$ $\text{mod }p_{1}$
which means that $q_{1}$ is a companion of $p_{1}$ with respect to $(p_{1},\widehat{p_{2}},p_{3})$. ∎
The following proposition gives three ways to think about the rational triangles studied in this paper.
Proposition 2.7.
Let $\Delta\subset\mathbb{R}^{2}$ be a triangle and consider a Markov triple $(p_{1},p_{2},p_{3})$. The following statements are equivalent.
  1. 1.
    $\Delta$ is a Markov triangle representing $(p_{1},p_{2},p_{3})$.
  2. 2.
    $\Delta$ is a pseudo-Markov triangle representing $(p_{1},p_{2},p_{3})$.
  3. 3.
    $\Delta$ is $\mathbb{R}^{2}\rtimes\operatorname{GL}(2,\mathbb{Z})$-congruent to a triangle having vertices
    $\Bigg\{(0,0),\Big(\frac{(p_{1}q_{1}-1)p_{2}}{p_{1}p_{3}},\frac{p_{1}p_{2}}{p_{3}}\Big),\Big(\frac{p_{3}}{p_{1}p_{2}},0\Big)\Bigg\}$ (4)
    for some companion number $q_{1}\equiv 3p_{3}p_{2}^{-1}$ mod $p_{1}$ of $p_{1}$.
Proof.
Let $v_{i,j}$ be the primitive vector along the $p_{i},p_{j}]-$edge of $\Delta$. Fix a companion number $q_{1}\equiv p_{3}p_{2}^{-1}$ mod $p_{1}$ of $p_{1}$, and let $\Delta_{0}$ be the triangle with $p_{1}$-, $p_{3}$-, and $p_{2}$-vertices given by (4).
[…]
$p_{3}^{2}=|a_{1}p_{2}^{2}-a_{2}p_{1}^{2}|=a_{1}p_{2}^{2}-a_{2}p_{1}^{2}.$ (5)
where we used the fact that $a_{1}p_{2}^{2}-a_{2}p_{1}^{2}=\det(v_{3,1},v_{3,2})>0$ because of our choice of orientation for $\Delta$: counterclockwise through $p_{1}$, then $p_{2}$, then $p_{3}$. From this and the Markov equation, we have
$a_{1}p_{2}^{2}=p_{3}^{2}+a_{2}p_{1}^{2}=3p_{1}p_{2}p_{3}-p_{1}^{2}-p_{2}^{2}+a_{2}p_{1}^{2}\equiv-p_{2}^{2}\qquad\text{mod }p_{1}$
and hence $a_{1}\equiv-1$ mod $p_{1}$ since $p_{1}$ and $p_{2}$ are coprime. In other terms, $a_{1}+1$ is divisible by $p_{1}$, allowing us to define the integer $q_{1}^{\prime}\vcentcolon=\frac{a_{1}+1}{p_{1}}$, for which $v_{1,3}=(p_{1}q_{1}^{\prime}-1,p_{1}^{2})$. We have
$q_{1}^{\prime}p_{2}^{2}=\frac{a_{1}p_{2}^{2}+p_{2}^{2}}{p_{1}}=\frac{3p_{1}p_{2}p_{3}-p_{1}^{2}+a_{2}p_{1}^{2}}{p_{1}}=3p_{2}p_{3}+p_{1}(a_{2}-1)$
so that $q_{1}^{\prime}p_{2}^{2}\equiv 3p_{2}p_{3}$ mod $p_{1}$. Multiplying through by $(p_{2}^{-1})^{2}$ yields $q_{1}^{\prime}\equiv 3p_{3}p_{2}^{-1}$ mod $p_{1}$. The coordinates of the $p_{3}$-vertex are then given by
$\Big(\frac{(p_{1}q_{1}^{\prime}-1)p_{2}}{p_{1}p_{3}},\frac{p_{1}p_{2}}{p_{3}}\Big)$
because the $p_{3}$-vertex belongs to the line $y=\frac{p_{1}^{2}}{p_{1}q_{1}^{\prime}-1}x$ spanned by $v_{1,3}$, and is at an affine distance $\frac{p_{2}}{p_{1}p_{3}}$ from the origin. Since $q_{1}\equiv q_{1}^{\prime}$ mod $p_{1}$, we are done by Remark 1.17.
[…]
$q_{1}p_{2}-3p_{3}\equiv 3p_{3}-3p_{3}\equiv 0\qquad\text{mod }p_{1}$
so that we can define the integer $q_{2}\vcentcolon=\frac{q_{1}p_{2}-3p_{3}}{p_{1}}$, which is in particular a companion number of $p_{2}$, namely $q_{2}\equiv-3p_{3}p_{1}^{-1}$ mod $p_{2}$. We first verify that $\Delta_{0}$ satisfies the edge property for being pseudo-Markov.
  • Dividing the vector from the $p_{2}$-vertex to the $p_{3}$-vertex by $\frac{p_{1}}{p_{2}p_{3}}$ yields $(p_{2}q_{1}+1,p_{2}^{2})$, which is primitive.
  • The vector $\frac{p_{1}p_{3}}{p_{2}}\big(\frac{(p_{1}q_{1}-1)p_{2}}{p_{1}p_{3}},\frac{p_{1}p_{2}}{p_{3}}\big)=(p_{1}q_{1}-1,p_{1}^{2})$ is primitive.
  • The affine length of the $p_{1},p_{2}]-$edge is of course $\frac{p_{3}}{p_{1}p_{2}}$.
We then verify the angle property.
  • At $p_{1}$: $v_{1,2}=(1,0)$ and $v_{1,3}=(p_{1}q_{1}-1,p_{1}^{2})$, so $|\det(v_{1,2},v_{1,3})|=p_{1}^{2}$.
  • At $p_{2}$: $v_{2,1}=(-1,0)$ and $v_{2,3}=(p_{2}q_{2}+1,p_{2}^{2})$, so $|\det(v_{2,1},v_{2,3})|=p_{2}^{2}$.
  • At $p_{3}$: $v_{3,1}=(1-p_{1}q_{1},-p_{1}^{2})$ and $v_{3,2}=(-p_{2}q_{2}-1,-p_{2}^{2})$, so
    $|\det(v_{3,1},v_{3,2})|$ $=|-p_{2}^{2}(1-p_{1}q_{1})+p_{1}^{2}(-p_{2}q_{2}-1)|$
    $=|-p_{1}^{2}-p_{2}^{2}+p_{2}^{2}p_{1}q_{1}-p_{1}^{2}p_{2}q_{2}|$
    $=|-p_{1}^{2}-p_{2}^{2}+p_{1}p_{2}(p_{2}q_{1}-p_{1}q_{2})|$
    $=|-p_{1}^{2}-p_{2}^{2}+3p_{1}p_{2}p_{3}|=|p_{3}^{2}|=p_{3}^{2}.$
[…]
Assume that (3) and (1) are equivalent for all (n1)(n-1)-st generation triples of the Markov tree. Consider the triangle $\widehat{\Delta}$ with $p_{1}$-, $\widehat{p_{2}}-$, and $p_{3}$-vertices
$\Bigg\{(0,0),\Big(\frac{(p_{1}q_{1}-1)p_{3}}{p_{1}\widehat{p_{2}}},\frac{p_{1}p_{3}}{\widehat{p_{2}}}\Big),\Big(\frac{\widehat{p_{2}}}{p_{1}p_{3}},0\Big)\Bigg\}.$
[…]
Claim.
$\widehat{\Delta}$ is the image of $\Delta_{0}$ under the geometric mutation $M_{p_{2}}$ at the $p_{2}$-vertex.
[…]
$y=\frac{p_{2}}{q_{2}}\big(x-\frac{p_{3}}{p_{1}p_{2}}\big)\qquad\text{and}\qquad y=\frac{p_{1}^{2}}{p_{1}q_{1}-1}x.$
Solving this system for $x$ yields
$x=\frac{p_{3}(p_{1}q_{1}-1)}{p_{1}\big(p_{1}(p_{2}q_{1}-p_{1}q_{2})-p_{2}\big)}.$
Using the expressions for $\widehat{p_{2}}$ and $q_{2}$ in terms of $p_{1},p_{2},p_{3},q_{1}$, we get
$x=\frac{p_{3}(p_{1}q_{1}-1)}{p_{1}(3p_{3}p_{1}-p_{2})}=\frac{p_{3}(p_{1}q_{1}-1)}{p_{1}\widehat{p_{2}}}\qquad\text{and}\qquad y=\frac{p_{1}^{2}}{p_{1}q_{1}-1}x=\frac{p_{1}p_{3}}{\widehat{p_{2}}}$
as claimed. We check that $M_{p_{2}}$ takes the $p_{3}$-vertex of $\Delta$ to the $p_{3}$-vertex of $\widehat{\Delta}$. The shear $\varphi_{v_{p_{2}}}$ is given by
$\varphi_{v_{p_{2}}}=\begin{pmatrix}1-x_{v_{p_{2}}}y_{v_{p_{2}}}&x_{v_{p_{2}}}^{2}\\ -y_{v_{p_{2}}}^{2}&1+x_{v_{p_{2}}}y_{v_{p_{2}}}\end{pmatrix}=\begin{pmatrix}1-q_{2}p_{2}&q_{2}^{2}\\ -p_{2}^{2}&1+q_{2}p_{2}\end{pmatrix}$. (6)
We compute
$M_{p_{2}}\begin{pmatrix}\frac{(p_{1}q_{1}-1)p_{2}}{p_{1}p_{3}}\\[6.0pt] \frac{p_{1}p_{2}}{p_{3}}\end{pmatrix}=\begin{pmatrix}\frac{q_{2}\big(p_{1}p_{2}(q_{2}p_{1}-p_{2}q_{1})+p_{2}^{2}\big)+p_{2}p_{1}q_{1}+q_{2}p_{3}^{2}-p_{2}}{p_{1}p_{3}}\\[6.0pt] \frac{p_{1}p_{2}^{3}(q_{2}p_{1}-q_{1}p_{2})+p_{2}^{2}(p_{1}^{2}+p_{2}^{2}+p_{3}^{2})}{p_{1}p_{2}p_{3}}\end{pmatrix}$.
Using the expressions for $\widehat{p_{2}}$ and $q_{2}$ again as well as the Markov equation, we simplify
$q_{2}\big(p_{1}p_{2}(q_{2}p_{1}-p_{2}q_{1})+p_{2}^{2}\big)$ $+p_{2}p_{1}q_{1}+q_{2}p_{3}^{2}-p_{2}$
$=q_{2}(-3p_{1}p_{2}p_{3}+p_{2}^{2})+p_{2}p_{1}q_{1}+q_{2}p_{3}^{2}-p_{2}$
$=q_{2}(-p_{1}^{2}-p_{3}^{2})+p_{2}p_{1}q_{1}+q_{2}p_{3}^{2}-p_{2}$
$=p_{1}(q_{1}p_{2}-q_{2}p_{1})-p_{2}=3p_{1}p_{3}-p_{2}=\widehat{p_{2}}$
and
$p_{1}p_{2}(q_{2}p_{1}-q_{1}p_{2})+p_{1}^{2}+p_{2}^{2}+p_{3}^{2}=-3p_{1}p_{2}p_{3}+p_{1}^{2}+p_{2}^{2}+p_{3}^{2}=0.$
Hence we have
$M_{p_{2}}\begin{pmatrix}\frac{(p_{1}q_{1}-1)p_{2}}{p_{1}p_{3}}\\[6.0pt] \frac{p_{1}p_{2}}{p_{3}}\end{pmatrix}=\begin{pmatrix}\frac{\widehat{p_{2}}}{p_{1}p_{3}}\\[6.0pt] 0\end{pmatrix}$
as desired. ∎
We observe that the geometric mutations of a Markov triangle preserve its integral barycentre. More precisely,
Lemma 2.8.
Let $\Delta$ be a Markov triangle representing $(p_{1},p_{2},p_{3})$ and $\beta$ the barycentre of $\Delta$. Then $\beta$ is also the barycentre of $M_{p_{i},\Delta}(\Delta)$ for $i\in\mathbb{Z}_{3}$.
Proof.
Let $\ell_{i}$ be the integral bisector at the $p_{i}$-vertex of $\Delta$. Without loss of generality, we fix the orientation of $\Delta$ to be counter-clockwise through $p_{1},p_{2},p_{3}$, and restrict our attention to the geometric mutation $M_{p_{2}}\vcentcolon=M_{p_{2},\Delta}$ at $p_{2}$. Thus $\widehat{\Delta}\vcentcolon=M_{p_{2}}(\Delta)$ is the triangle with vertices $p_{1}$, $M_{p_{2}}(p_{3})$, and $\widehat{p_{2}}\vcentcolon=\ell_{2}\cap[p_{1},p_{3}]$. We denote the edges of $\widehat{\Delta}$ by $\widehat{e_{1,2}}\vcentcolon=[p_{1},\widehat{p_{2}}$], $\widehat{e_{1,3}}\vcentcolon=[p_{1},M_{p_{2}}(p_{3})]$, and $\widehat{e_{2,3}}\vcentcolon=[\widehat{p_{2}},M_{p_{2}}(p_{3})]$ and let $\widehat{v_{i,j}}$ be the primitive integral vector along $\widehat{e_{i,j}}$. We observe that $\widehat{v_{1,3}}=v_{1,2}$ and $\widehat{v_{1,2}}=v_{1,3}$, so that the integral bisector $\widehat{\ell_{1}}$ at the $p_{1}$-vertex of $\widehat{\Delta}$ is still $\ell_{1}$. On the other hand, we claim that the integral bisector $\widehat{\ell_{2}}$ at the $\widehat{p_{2}}-$vertex is $\ell_{2}$, because the points lying on $\ell_{2}$ are at an equal affine distance from both edges $\widehat{e_{1,2}}$ and $\widehat{e_{2,3}}$. This property is trivially satisfied by the $\widehat{p_{2}}-$vertex, so we only need to find one other point of $\ell_{2}$ satisfying it. Consider the $p_{2}$-vertex of $\Delta$. The affine distance from $p_{2}$ to $\widehat{e_{2,3}}$ is given by the absolute value of the determinant of $v\vcentcolon=M_{p_{2}}(p_{3})-p_{2}$ with $\widehat{v_{2,3}}$. Since $v$ is the image of $p_{3}-p_{2}$ under a shear, both vector have the same affine length, equal to $\frac{p_{1}}{p_{2}p_{3}}$ by the edge property of $\Delta$ as a pseudo-Markov triangle in Definition 2.5. Thus we have
$\operatorname{dist}(p_{2},\widehat{e_{2,3}})=|\det(v,\widehat{v_{2,3}})|=\frac{p_{1}}{p_{2}p_{3}}|\det(v_{1,2},\widehat{v_{2,3}})|=\frac{p_{2}p_{3}}{p_{2}}$
where $|\det(v_{2,1},\widehat{v_{2,3}})|=p_{3}^{2}$ by the angle property of $\widehat{\Delta}$ as a pseudo-Markov triangle. Likewise the affine distance from $p_{2}$ to $\widehat{e_{1,2}}$ is given by
$\operatorname{dist}(p_{2},\widehat{e_{1,2}})=|\det(v,\widehat{v_{1,2}})|=\frac{p_{3}}{p_{1}p_{3}}|\det(v_{1,2},\widehat{v_{1,3}})|=\frac{p_{2}p_{3}}{p_{2}}$
We conclude that $\widehat{\ell_{2}}=\ell_{2}$. Therefore the integral barycentre $\widehat{\beta}$ of $\widehat{\Delta}$ is given by $\widehat{\beta}=\widehat{\ell_{1}}\cap\widehat{\ell_{2}}=\ell_{1}\cap\ell_{2}=\beta$. ∎
$p_{1}$$p_{2}$$p_{3}$$\ell_{1}$$\ell_{2}$$\beta$$\Delta$$\widehat{\Delta}$$\widehat{p_{2}}$$M_{p_{2}}(p_{3})$$\frac{p_{1}}{p_{2}p_{3}}$$\widehat{v_{2,3}}$$v_{1,3}$$v_{1,2}$
Figure 10: Preservation of the integral barycentre under geometric mutation
Remark 2.9.
It follows directly from Lemma 2.8 and its proof that the affine distance from $\beta$ to any edge of $\Delta$ is preserved by the geometric mutation of $\Delta$ at any vertex.
2.3 On the Ehrhart theory of rational Markov triangles
Proof of Proposition 1.5.
Let $P\subset\mathbb{R}^{2}$ be a polygon and $v$ an integral vector. Consider the half-shear $\varphi\vcentcolon=\varphi_{v}^{1/2}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ with respect to $v$, which preserves the integer lattice as mentioned in Remark 1.3. Then for any point $x\in P$ and integer dilation factor $t\in\mathbb{N}\cup\{0\}$, $tx\in\mathbb{Z}^{2}$ if and only if $t\varphi(x)=\varphi(tx)\in\mathbb{Z}^{2}$, so that $\varphi$ defines a bijection between the integer lattice points in the $t$-dilate of $P$ and in the $t$-dilate of $\varphi(P)$, for each $t\in\mathbb{N}\cup\{0\}$. Since $P$ has finitely many integer lattice points, we get
$L_{P}(t)=\#(tP)\cap\mathbb{Z}^{2}=\#\big(t\varphi(P)\cap\mathbb{Z}^{2}\big)=L_{\varphi(P)}(t).\qed$
Proof of Proposition 1.20.
Let $(p_{1},p_{2},p_{3})$ be a Markov triple and consider a representative Markov triangle $\Delta^{\beta}$ in standard $p_{1}$-barycentric position. For each $i=1,2,3$, the integral barycentre $\beta$ of $\Delta$ lies on the integral bisector $\ell_{i}$ of the $p_{i}$-vertex, so that the position vector $p_{i}=p_{i}-\beta$ is parallel to $\ell_{i}$. It follows that the half-shear $\varphi_{v_{i}}^{1/2}$ with respect to the primitive vector $v_{i}$ along $\ell_{i}$ is invariant under translation along $p_{i}$, and hence the geometric mutation $M_{p_{i}}=T_{p_{i}}\circ\varphi_{v_{i}}^{1/2}\circ T_{-p_{i}}=\varphi_{v_{i}}^{1/2}$ is the half-shear itself, so that $M_{p_{i}}(\Delta^{\beta})$ and $\Delta^{\beta}$ are Ehrhart equivalent by Proposition 1.5. ∎
For the proof of Proposition 1.21, we need the following two results on the effect of geometric mutations on Markov triangles in standard $p$-position.
Lemma 2.10.
Let $(p_{1},p_{2},p_{3})$ be a Markov triple and $\Delta$ a representing triangle in standard $p_{1}$-position. Let $\widehat{\Delta}$ be the image of $\Delta$ under its geometric mutation at any vertex. Then $p_{1}\Delta$ and $p_{1}\widehat{\Delta}$ are Ehrhart equivalent.
Proof.
In the following we assume that the $p_{2}$-vertex lies on the $x$-axis. The result is trivial for the geometric mutation of the $t$-dilate of $\Delta$ at $tp_{1}$, since $tp_{1}$ is the origin and hence $M_{tp_{1},t\Delta}=M_{(0,0),t\Delta}$ is the half-shear along $v_{1}$, and in particular preserves the integer lattice. Thus $M_{p_{1},\Delta}(\Delta)$ and $\Delta$ are Ehrhart equivalent.
We next consider the geometric mutation of the $t$-dilate of $\Delta$ at $tp_{2}$, defined as
$M_{tp_{2},t\Delta}=T_{(\frac{tp_{3}}{p_{1}p_{2}},0)}\circ\varphi_{v_{p_{2}}}^{1/2}\circ T_{(\frac{-tp_{3}}{p_{1}p_{2}},0)}$
as described in the proof of Proposition 2.7, where the half-shear $\varphi_{v_{p_{2}}}^{1/2}$ is given on $H_{v_{p_{2}}}^{+}$ in matrix form by (6). Take an integer point $(x,y)\in\mathbb{Z}^{2}$. If $(x,y)$ lies to the left of the integral bisector $\ell(tp_{2},v_{p_{2}})$ at the $tp_{2}$-vertex of $t\Delta$, then $M_{tp_{2},t\Delta}(x,y)=(x,y)$. Otherwise,
$M_{tp_{2},t\Delta}(x,y)$ $=\begin{pmatrix}\frac{tp_{3}}{p_{1}p_{2}}\\[6.0pt] 0\end{pmatrix}+\begin{pmatrix}1-p_{2}q_{2}&q_{2}^{2}\\[6.0pt] -p_{2}^{2}&1+p_{2}q_{2}\end{pmatrix}\begin{pmatrix}x-\frac{tp_{3}}{p_{1}p_{2}}\\[6.0pt] y\end{pmatrix}$
$=\begin{pmatrix}x(1-p_{2}q_{2})+y{q_{2}}^{2}+q_{2}p_{3}\frac{t}{p_{1}}\\[6.0pt] x(-p_{2}^{2})+y(1+p_{2}q_{2})+p_{2}p_{3}\frac{t}{p_{1}}\end{pmatrix}$
$=\begin{pmatrix}k_{1}x+k_{1}^{\prime}y+k_{1}^{\prime\prime}\frac{t}{p_{1}}\\[6.0pt] k_{2}x+k_{2}^{\prime}y+k_{2}^{\prime\prime}\frac{t}{p_{1}}\end{pmatrix}$
for some integers $k_{1},k_{1}^{\prime},k_{1}^{\prime\prime},k_{2},k_{2}^{\prime},k_{2}^{\prime\prime}\in\mathbb{Z}$. Hence $M_{tp_{2},t\Delta}(\mathbb{Z}^{2})\subset\mathbb{Z}^{2}$ when $p_{1}$ divides $t$; the sublattice $M_{tp_{2},t\Delta}(\mathbb{Z}^{2})$ of $\mathbb{Z}^{2}$ has index $\det(M_{tp_{2},t\Delta})=1$, so that $M_{tp_{2},t\Delta}(\mathbb{Z}^{2})=\mathbb{Z}^{2}$. Thus $p_{1}\Delta$ and $p_{1}\widehat{\Delta}$ are Ehrhart equivalent.
Finally we consider the geometric mutation of $t\Delta$ at $tp_{3}$. Applying the integral map $f\vcentcolon=\psi_{p_{1}}^{p_{2},p_{3}}$ defined in Remark 1.17 to $\Delta$ yields the Markov triangle $f(\Delta)$ representing $(p_{1},p_{2},p_{3})$ in standard $p_{1}$-position with the $p_{3}$-index lying on the $x$-axis. We repeat the argument of the mutation at $tp_{2}$ after swapping $p_{2}$ and $p_{3}$ and replacing $q_{2}$ by $q_{3}\vcentcolon=\frac{-q_{1}p_{3}-3p_{2}}{p_{1}}$, and conclude that $M_{tp_{3},tf(\Delta)}(\mathbb{Z}^{2})=\mathbb{Z}^{2}$ when $t$ divides $p_{1}$.
Lemma 2.11.
Let $m_{n}$ be a Markov number occurring as the largest element in a triple $\mathfrak{m}$ of the $n$-th generation of the Markov tree, and let $\Delta$ be a Markov triangle representing $\mathfrak{m}$ in standard $m_{n}$-position. There exist n1n-1 Markov numbers $m_{1},\cdots,m_{n-1}$ less than $m_{n}$ such that the $m_{1}\cdots m_{n}$-dilate of $\Delta$ is Ehrhart equivalent to the $m_{1}\cdots m_{n}$-dilation of a triangle $\Delta_{(1,1,1)}$ representing $(1,1,1)$ in standard $1$-position.
Proof.
[…]
Figure 11: Geometric mutation of $\Delta$ at $m_{n}$
Proof of Proposition 1.21.
Let $\Delta$ be a $(p_{1},p_{2},p_{3})$-triangle in standard $p_{1}$-position with some companion number $q_{1}\equiv 3p_{3}p_{2}^{-2}$ mod $p_{1}$ of $p_{1}$.
Assertion 1.
The period of $L_{\Delta}$ is a divisor of $3p_{1}$.
Proof of Assertion 1. It suffices to show that the $3p_{1}$-dilate of $\Delta$ is Ehrhart equivalent to the $3p_{1}$-dilation of a triangle representing $(p_{1},p_{2},p_{3})$ in standard $p_{1}$-barycentric position, which is pseudo-integral by Proposition 1.20. To this end we compute the coordinates of the integral barycentre $\beta$ of $\Delta$. By the computation in the proof of Proposition 2.7, the primitive vector $v_{1}$ along the integral bisector $\ell_{1}$ at the $p_{1}$-vertex is a positive multiple of $(q_{1},p_{1})$. But $q_{1}$ is relatively prime to $p_{1}$ by Lemma 2.6, so that $v_{1}=(q_{1},p_{1})$. Thus $\ell_{1}$ is defined by the equation $y=\frac{p_{1}}{q_{1}}x$. By Remark 2.9, the affine distance from $\beta$ to each of the edges of $\Delta$ is preserved by geometric mutations, and is therefore equal to $1/3$. Hence the affine distance from $\beta$ to the $p_{1},p_{2}]-$edge with primitive vector $(1,0)$ is
$\frac{1}{3}=|\det\big((1,0),\beta\big)|=\beta_{y}.$
Since $\beta$ lies on $\ell_{1}$, we conclude that $\beta_{x}=\frac{q_{1}}{3p_{1}}$. Thus the $t$-dilate of $\Delta$ is the translation of the $t$-dilation of a $(p_{1},p_{2},p_{3})$-triangle $\Delta^{\prime}$ in standard $p_{1}$-barycentric position along the vector $(\frac{q_{1}t}{3p_{1}},\frac{t}{3})$. This translation preserves the lattice whenever $3p_{1}$ divides $t$, so $t\Delta$ and $t\Delta^{\prime}$ have the same number of integer lattice points if $3p_{1}\mid t$. In other terms, $3p_{1}\Delta$ and $3p_{1}\Delta^{\prime}$ are Ehrhart equivalent.
Assertion 2.
The period of $L_{\Delta}$ is a divisor of $kp_{1}$, where $k$ is a product of finitely many Markov numbers which are smaller than $p_{1}$.
Proof of Assertion 2. Let $(p_{2}^{\prime},p_{1},p_{3}^{\prime})$ be the root of the bivalent tree $\Lambda(p_{2}^{\prime},p_{1},p_{3}^{\prime})$ to which $(p_{1},p_{2},p_{3})$ belongs, and let $n$ be the generation of $(p_{2}^{\prime},p_{1},p_{3}^{\prime})$ and $m$ the generation of $(p_{1},p_{2},p_{3})$ in the Markov tree. Applying mnm-n geometric mutations to $\Delta$ (corresponding to decreasing mutations of Markov triples) yields a Markov triangle $\Delta_{0}$ representing $(p_{2}^{\prime},p_{1},p_{3}^{\prime})$ in standard $p_{1}$-position. By Lemma 2.10, $p_{1}\Delta_{0}$ and $p_{1}\Delta$ are Ehrhart equivalent. By Lemma 2.11, $k\Delta_{0}$ is Ehrhart equivalent to $k\Delta_{(1,1,1)}$ where $k$ is the product of n1n-1 Markov numbers smaller than $p_{1}$; it follows that $kp_{1}\Delta$ is Ehrhart equivalent to $kp_{1}\Delta_{(1,1,1)}$, which is integral. Therefore the period of $L_{\Delta}$ divides $kp_{1}$.
By Assertions 1 and 2, the period of $L_{\Delta}$ is a common divisor of $3p_{1}$ and $kp_{1}$, and hence divides $\operatorname{gcd}(3p_{1},kp_{1})$. Since every Markov number is relatively prime to $3$ and $k$ is a product of Markov numbers, it follows that $\operatorname{gcd}(3p_{1},kp_{1})=p_{1}$, as claimed. ∎
Proof of Corollary 1.22.
By Proposition 1.21, the period of such a triangle divides $2$. The claim follows by observing that the triangle $\Delta$ with vertices $(0,0)$, $(1/2,2)$, and $(1/2,0)$ is not pseudo-integral, as seen by computing $L_{\Delta}(t)$ for $t=1,2,3$. It follows that the period cannot be 11 and must therefore be 22. ∎
2.4 On the Ehrhart theory of irrational Markov triangles
We begin by asserting the Hausdorff convergence of the sequences of Markov triangles in standard $p_{1}$-position. Let $e_{\infty}$ be the $b_{\infty},c_{\infty}]-$edge of $\Delta_{\infty}^{a}$.
Proof of Proposition 1.24.
Let $C\vcentcolon=\operatorname{Cone}(v_{a}^{(1)},v_{a}^{(2)})$ be the cone generated by the primitive vectors emanating from the $a$-vertex of $\Delta_{\infty}^{a}$ (or of $\Delta_{n}^{a}$ for any $n\in\mathbb{N}$, since $v_{a}^{(1)},v_{a}^{(2)}$ are preserved by the geometric mutations involved in the definition of $\Delta_{n}^{a}$). For any $\varepsilon>0$, let $U_{\varepsilon}$ be the $\varepsilon$-neighbourhood of $e_{\infty}$ inside $C$, i.e. the set of all points in $C$ lying at a Euclidean distance smaller than $\varepsilon$ from $e_{\infty}$. The limits (3) mean that the sequence of $b_{n}$-vertices of $\Delta_{n}^{a}$ converges to the $b_{\infty}$-vertex of $\Delta^{a}_{\infty}$, and similarly the sequence of $c_{n}$-vertices of $\Delta_{n}^{a}$ converges to the $c_{\infty}$-vertex of $\Delta^{a}_{\infty}$. In particular for our choice of $\varepsilon$ there exists $N\in\mathbb{N}$ for which the $b_{n}$- resp. $c_{n}$-vertex lies within a distance of $\varepsilon$ from the $b_{\infty}$- resp. $c_{\infty}$-vertex and hence the $b_{n},c_{n}]-$edge lies entirely inside $U_{\varepsilon}$ for $n\geq N$. Since the complements of $U_{\varepsilon}$ in our triangles agree for $n$ sufficiently large, namely $\Delta_{\infty}^{a}\setminus U_{\varepsilon}=\Delta_{n}^{a}\setminus U_{\varepsilon}$ for $n\geq N$, it follows that the symmetric difference $\Delta_{n}^{a}\ominus\Delta_{\infty}^{a}\subset U_{\varepsilon}$ is entirely contained in $U_{\varepsilon}$, which is arbitrarily small. ∎
[…]
Figure 12: Representation of a neighbourhood of the symmetric difference $\Delta_{n}^{a}\ominus\Delta_{\infty}^{a}$
Remark 2.12.
Hausdorff convergence does not imply Ehrhart equivalence: given a sequence $T_{n}$ of triangles Hausdorff converging to some triangle $T_{\infty}$, a lattice point may lie in the exterior of $T_{n}$ for all $n\in\mathbb{N}$ and on the boundary of $T_{\infty}$, as shown by the following example. Consider the sequence $(T_{n})$ where $T_{n}$ is the triangle with vertices
$\Bigg\{\Big(0,3\frac{\operatorname{Fib}_{2n+1}}{\operatorname{Fib}_{2n+3}}\Big),\Big(3\frac{\operatorname{Fib}_{2n+3}}{\operatorname{Fib}_{2n+1}},0\Big),\Big(3\frac{\operatorname{Fib}_{2n+3}}{\operatorname{Fib}_{2n+1}},3\frac{\operatorname{Fib}_{2n+1}}{\operatorname{Fib}_{2n+3}}\Big)\Bigg\}$
which Hausdorff converges to the triangle $T_{\infty}$ with vertices
$\Big\{(0,\frac{3}{\tau^{2}}),(3\tau^{2},0),(3\tau^{2},\frac{3}{\tau^{2}})\Big\}.$
Nevertheless, for all $n\in\mathbb{N}$, $T_{\infty}$ and $T_{n}$ are not Ehrhart equivalent, since $T_{n}$ contains the integer lattice points $(k,1)$ for $k=2,\cdots,7$, while $T_{\infty}$ contains in addition the point $(1,1)$.
Remark 2.12 tells us that to understand the Ehrhart function of $\Delta_{\infty}^{a}$, we must investigate its boundary integer lattice points. The edge $e_{\infty}$ spans the affine line
$D_{\infty}:a^{2}x+\big(a^{2}\lambda(a)^{2}-(aq-1)\big)y=a^{2}\lambda(a).$
Since $9a^{2}-4$ is not a perfect square, $\sqrt{9-4/a^{2}}$ is irrational. Therefore the slope of $D_{\infty}$ is irrational and hence $D_{\infty}$ contains exactly one rational point. We identify it in the following
Lemma 2.13.
The only rational point of the affine line $D_{\infty}$ is $\beta\vcentcolon=(\frac{q}{3a},\frac{1}{3})$, namely the integral barycentre of $\Delta_{n}$ for all $n\in\mathbb{N}$.
Proof.
Direct computation shows that the equation
$a^{2}x+\big(a^{2}\lambda(a)^{2}-(aq-1)\big)y=a^{2}\lambda(a)$
is satisfied by $(\frac{q}{3a},\frac{1}{3})$. ∎
Lemma 2.14.
For each $t\in\mathbb{N}$, $te_{\infty}$ has a neighbourhood $U_{t}$ with $U_{t}\cap\mathbb{Z}^{2}\setminus\{t\beta\}=\varnothing$.
Proof.
Fix $t\in\mathbb{N}$. Let $d_{(t)}:\mathbb{R}^{2}\rightarrow\mathbb{R}_{\geq 0}$ be the continuous function mapping each point of the plane to its Euclidean distance from $tD_{\infty}$. $d_{(t)}$ clearly attains its minimum $\varepsilon\vcentcolon=\min\{d_{(t)}(x,y)\mid(x,y)\in C\cap\mathbb{Z}^{2}\setminus\beta\}$ over the integer lattice points in the cone $C$ punctured at $\beta$. It thus follows from Lemma 2.13 that $\varepsilon$ is positive. We define $U_{t}\vcentcolon=\{(x,y)\in C\setminus\beta\mid d_{(t)}(x,y)<\varepsilon\}$ as the $\varepsilon$-tubular neighbourhood of $te_{\infty}$ in $C$. By construction $U_{t}\setminus\{t\beta\}$ contains no integer lattice point. ∎
Lemma 2.14 allows us to identify the integer lattice points lying in $t\Delta_{\infty}^{a}$:
Proposition 2.15.
For each $t\in\mathbb{N}$, there is an $N_{t}\in\mathbb{N}$ such that $t\Delta_{\infty}^{a}\cap\mathbb{Z}^{2}=t\Delta_{n}^{a}\cap\mathbb{Z}^{2}$ for all $n\geq N_{t}$.
Proof.
Fix $t\in\mathbb{N}$, and consider the neighbourhood $U_{t}$ of $e_{\infty}$ given by Lemma 2.14 so that $U_{t}\cap\mathbb{Z}^{2}\setminus\{t\beta\}=\varnothing$. As in the proof of Proposition 1.24, we see that the sequence of $t$-dilated Markov triangles $(t\Delta_{n}^{a})$ that this sequence Hausdorff converges to the $t$-dilated limiting triangle $t\Delta_{\infty}^{a}$, and therefore there exists $N_{t}\in\mathbb{N}$ such that the symmetric difference $t\Delta_{n}^{a}\ominus t\Delta_{\infty}^{a}$ is entirely contained in $U_{t}$. We note that the point $t\beta$ belongs to both $t\Delta_{n}^{a}$ as its integral barycentre and to $t\Delta_{\infty}^{a}$ by Lemma 2.13. Thus the complements $t\Delta_{n}^{a}\setminus(U_{t}\setminus\{t\beta\})$ and $t\Delta_{\infty}^{a}\setminus(U_{t}\setminus\{t\beta\})$ of the punctured neighbourhood are equal and hence contain the same integer lattice points. On the other hand both intersections $t\Delta_{n}^{a}\cap(U_{t}\setminus\{t\beta\})$ and $t\Delta_{\infty}^{a}\cap(U_{t}\setminus\{t\beta\})$ contain no lattice point by definition of $U_{t}$. ∎
Proof of Theorem 1.25.
By Proposition 2.15, for any $t\in\mathbb{N}$, $at\Delta_{\infty}^{a}$ and $at\Delta_{N_{t}}^{a}$ have the same integer lattice points for some $N_{t}\in\mathbb{N}$, i.e. $L_{a\Delta_{N_{t}}^{a}}(t)=L_{a\Delta_{\infty}^{a}}(t)$. Since all $a$-dilations of the triangles $\{\Delta_{n}^{a}\}_{n\in\mathbb{N}}$ are Ehrhart equivalent by the proof of Proposition 1.21, $L_{a\Delta_{n}^{a}}(t)$ does not depend on $n$, so that $L_{a\Delta_{\infty}^{a}}(t)=L_{a\Delta_{n}^{a}}(t)$ for all $n\in\mathbb{N}$. Since this holds for all $t$, we conclude that $L_{a\Delta_{\infty}^{a}}=L_{a\Delta_{1}^{a}}$. ∎
Proof of Theorem 1.26.
We denote the $b_{\infty},c_{\infty}]-$edge of $\Delta_{\infty}^{\beta}$ by $e_{\infty}^{\beta}$, which spans the line $D_{\infty}^{\beta}$. By Lemma 2.13, the only rational point of $D_{\infty}^{\beta}$ is the origin $\beta=(0,0)$. Arguing similarly as in the proof of Lemma 2.14, we construct a neighbourhood $U_{t}$ of $te_{\infty}^{\beta}$ whose unique integer lattice point is $(0,0)$. Again arguing similarly as in the proof of Proposition 2.15 implies, for each $t\in\mathbb{N}$, the existence of $N_{t}\in\mathbb{N}$ such that $t\Delta_{\infty}^{\beta}\cap\mathbb{Z}^{2}=t\Delta_{n}^{\beta}\cap\mathbb{Z}^{2}$ for all $n\geq N_{t}$. Finally the proof of Theorem 1.25 directly applies in this case as well and we are done. ∎